For the most recent year available, the mean annual cost to attend a private university in the United States was $20,132. Assume the distribution of annual costs follows the normal probability distribution and the standard deviation is $4,450.


Ninety percent of all students at private universities pay less than what amount? (Round z value to 2 decimal places and your final answer to the nearest whole number.)

Amount $
1
Expert's answer
2013-03-05T09:54:34-0500
We have normal distribution with mean $20,132 and standard deviation - $4,450:
m = 20,132
sigma = 4,450
And we know:
P ( x < mean + z*sigma ) = 0.9
or& percentage of values expected to lie in and outside the symmetric interval& mean +- z*sigma = 80 %
From the tables: z = 1.281552 = 1.28
Finally,
z*sigma = 1.28*4450 = 5696
Therefore, 90% of all students at private universities pay less than :
20,132+5696=25828
Answer: z=1.28,& Amount& $25828